OpenStudy (anonymous):

Can someone please help me? I dont know how to work this problem! Perform the operation and leave the result in trigonometric form. 4(cos40deg+isin40deg) ----------------------- 3(cos160deg+isin160deg)

OpenStudy (anonymous):

Thats divided btw.

zepdrix (zepdrix):

\[\Large\bf\sf 3\left[\cos(4\cdot40)+i \sin(4\cdot40)\right]\]De'Moire's Theorem (I hope I spelled that right) allows us to write it like this:\[\Large\bf\sf 3\left[\cos40+i \sin40\right]^4\]Bringing the 4 outside as an exponent. Does that identity look familiar?

OpenStudy (anonymous):

The answer I got was 4(cos 40 + isin 40)^3 Is that correct? Or is your answer correct?

zepdrix (zepdrix):

Hmmm,\[\Large\bf\sf \cos n \theta+ i \sin n \theta \quad\ne\quad (\cos n \theta+i \sin n \theta)^n\]You've got to take the n out of our trig functions. I have a feeling that's what you did.. hmm Darn I gotta go to work D: sorry

OpenStudy (anonymous):

Alright let me see. You have to take out the cos and the isin? I dont understand. Can you show me? I have other problems like this one to solve.

OpenStudy (anonymous):

i take it you are dividing right?

OpenStudy (anonymous):

the number out front stays \(\frac{4}{3}\) then subtract the angles that is all

OpenStudy (anonymous):

there are lots of ways to write it but your basic job is to compute \(40-160\)

OpenStudy (anonymous):

I am still lost. I'm sorry. I dont like these problems, lol